a spy satellite is in circular orbit around earth. it makes one revolution in 6.00 hr, how high above the earth's surface is the satellite?
1
Expert's answer
2019-01-28T15:58:52-0500

Let's first find the distance of the satellite from the center of the Earth from the Kepler's Third Law:

"\frac{T^2}{R^3} = \frac{4\pi^2}{GM_E},"

here,

"T"

is the period that the satellite needs to make one revolution,

"R"

is the distance of the satellite from the center of the Earth,

"G"

is the universal gravitational constant and

"M_E"

is the mass of the Earth.

Then, from this equation we can find the distance

of the satellite from the center of the Earth:

"R = \sqrt[3]{\frac{GMT^2}{4\pi^2}}."

Let's substitute the numbers:

"R = \sqrt[3]{\frac{6.67 \cdot 10^{-11} \frac{Nm^2}{kg^2} \cdot 5.98 \cdot 10^{24} kg \cdot (6.0 hr \cdot \frac{3600 s}{1 hr})^2}{4\pi^2}} = 1.67 \cdot 10^7 m."

We can find the height of the satellite above the Earth's surface from the formula:

"R = R_{Earth} + h,"

here,

"R_{Earth}"

is the radius of the Earth,

"h"

is the height of the satellite above the Earth's surface.

Finally, we get:

"h = R - R_{Earth} = 1.67 \cdot 10^7 m - 6.37 \cdot 10^6 m = 1.03 \cdot 10^7 m."

Answer:

"h = 1.03 \cdot 10^7 m"

.